How long will each question take to answer if there are 80 questions in a 4 hour long test?
Found 2 solutions by Theo, ikeoddy:
Answer by Theo(3458)

4 hours * 60 = 240 minutes.
240 / 80 = 3 minutes for each question on the average.
some will take more.
some will take less.
the overall average, however, needs to be less than or equal to 3 minutes per question if you expect to answer them all.

Answer by ikeoddy(3)

Assume that each question takes the same amount of time to answer. Let this time be x minutes.

80 questions can be answered in 4 hours = 4*60 minutes = 240 minutes
1 question can be answered in x minutes

If we utilize proportions,
240/80 = x/1

Thus, x = 3.

Hence, it will take 3 minutes to answer each question.